OpenStudy (anonymous):

Write the equation in standard form: 9(x+2)^2+5y^2=45

OpenStudy (ahsome):

\[9(x+2)^2+5y^2=45\] Right?

OpenStudy (anonymous):

Yes.

OpenStudy (ahsome):

\[9(x+2)^2+5y^2=45\] \[9(x^2+4x+4)+5y^2=45\] \[9x^2+36x+36+5y^2=45\] \[9x^2+36x+5y^2=9\] \[5y^2=-9x^2-36x+9\]

OpenStudy (anonymous):

How would I go about graphing it?

OpenStudy (anonymous):

WITH PAPER.